# Artificial Intelligence LLM Large Language Model

### v2

@@ Abstract.
As the use of Large Language Model (LLM) chatbots by students and researchers becomes more prevalent, universities are pressed to develop AI strategies. One strategy that many universities pursue is to customize pre-trained LLM as-a-service (LLMaaS). While most studies on LLMaaS chatbots prioritize technical adaptations, we focus on psychological effects of user-salient customizations, such as interface changes. We assume that such customizations influence users’ perception of the system and are therefore important in guiding safe and appropriate use. In a field study, we examine how students and employees ($N=526$) at a German university perceive and use their institution’s customized LLMaaS chatbot compared to ChatGPT. Participants using both systems ($n=116$) reported greater trust, higher perceived privacy and less experienced hallucinations with their university’s customized LLMaaS chatbot in contrast to ChatGPT. We discuss theoretical implications for research on calibrated trust, and offer guidance on the design and deployment of LLMaaS chatbots.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Natural language interfacesccs: Security and privacy Privacy protections
@@ 1. Introduction
Large Language Model (LLM) chatbots are becoming increasingly prevalent in educational settings. Current research shows high adoption rates of commercial off-the-shelf LLM chatbots (e. g., ChatGPT) among students and university staff (Zhang et al., 2024; Zhou et al., 2023; von Garrel and Mayer, 2023). Consequently, universities face mounting pressure to develop strategies for AI integration (Kapania et al., 2025). This involves multiple decisions, such as whether and how to provide employees with access to LLMs. To guide the informed and secure use of such systems, many organizations, including universities, corporations, and government agencies, are exploring the deployment of LLM as-a-Service (LLMaaS) as a feasible and cost-effective way to adopt AI in-house (La Malfa et al., 2023; Chkirbene et al., 2024). These services can form the basis for internal LLMaaS-based chatbots in organizations. The adoption of such services is driven by several practical considerations, including the opportunity to offer access to all university members while maintaining institutional oversight and alignment with organizational policies (Paavola, 2024). One of the main benefits of AI-as-a-Service (AIaaS) solutions, of which LLMaaS is a specific instance, is the flexibility they offer in customizing various aspects (Diaferia et al., 2022), such as fine-tuning models or displaying user interfaces in corporate design.
@@
Customizations of LLMaaS solutions play a critical role, as they can enhance model performance within specific organizational contexts and strengthen security and privacy measures (Sundberg and Holmström, 2023; Lewicki et al., 2023). However, even choices that do not impact model performance or security, such as branding the user interface, could lead to significant changes in how the system is perceived and used. Psychological research suggests that people draw inferences about a system’s trustworthiness based on indicators available to them, also referred to as cues (de Visser et al., 2014), a mechanism formalized in the Trustworthiness Assessment Model (TrAM) (Schlicker et al., 2025). Real-world incidents show the importance of understanding user perception of LLM chatbots: cases like the data leak by Samsung employees via ChatGPT illustrate how overly trusting a system or having false privacy expectations can lead to the unintended disclosure of sensitive information (Mauran, 2023). Such incidents underscore the role of trust calibration, meaning the degree to which users’ perceived trustworthiness of a system aligns with its actual properties and limitations (Lee and See, 2004; Parasuraman and Riley, 1997). While recent studies have focused on how to intentionally improve calibrated trust and reliance on LLMs through interventions (Bo et al., 2025; Kim et al., 2024; Ma et al., 2025a; Vasconcelos et al., 2025), we lack research on how customization choices of real-world LLM chatbots impact the calibration of user perceptions. In particular, research is lacking on customization choices that are salient to users (e. g., in the form of cues perceivable in the user interface) but have no impact on aspects such as the model, training data, privacy safeguards, or system infrastructure, which have been the primary focus of prior work (Diaferia et al., 2022; Ali et al., 2019; Tönjes et al., 2024).
@@
In this work, we examine if cues – emerging from different customization choices in LLMaaS chatbots – influence users’ perceptions and usage. We draw on the concept of trust calibration (Lee and See, 2004; Parasuraman and Riley, 1997) and the TrAM (Schlicker et al., 2025) to empirically investigate the alignment (or misalignment) between users’ perception of a customized LLMaaS and the system’s actual characteristics, thereby answering the following research question:
RQ: How do customizations of LLMaaS chatbots impact end users’ trust, perception, and usage of LLM chatbots in professional and academic contexts compared to commercial AI alternatives?
@@
To examine this research question, we conducted a survey at a major German university. A diverse sample of students and university employees (N = 526) answered questions about their experiences with the university’s customized LLMaaS chatbot and with ChatGPT (the most widely used LLM chatbot among German students (von Garrel and Mayer, 2025)). We measured trust, privacy concerns, cautious behavior toward LLM hallucinations, experienced LLM hallucinations, and sustainability-aware AI usage. Our main analyses focus on a within-subject subsample of participants (n = 116) who regularly use both systems, allowing direct comparison of user perceptions and behaviors across the two platforms.
@@
    We show that participants indicate greater trust, higher perceived privacy and less experienced LLM hallucinations with their university’s customized LLMaaS chatbot compared to ChatGPT.
@@
    Drawing on the TrAM framework (Schlicker et al., 2025), we demonstrate that customizations can serve as cues influencing users’ perceptions and behavior toward LLMaaS chatbots.
@@ 2.1. LLM (as-a-Service) Chatbots at Universities
LLM chatbots, such as ChatGPT, are widely used in academic settings, with significant adoption among university students (von Garrel and Mayer, 2023; Amani et al., 2023) and researchers (Kapania et al., 2025; Syed et al., 2024; Abdelhafiz et al., 2024). Thus, universities are caught between enabling students and staff to access and use AI in an informed way to leverage the benefits while also protecting data, privacy, rights, and ethical principles. Introducing “their own” LLMs poses a considerable challenge for universities, due to a lack of internal technical abilities and knowledge about the deployment of AI and the high costs of developing and maintaining a sufficient IT infrastructure (Lins et al., 2021).
@@
LLM as-a-Service (LLMaaS) provide a practical solution by offering easy access to pre-trained models via cloud providers (Wang et al., 2024; Syed et al., 2025). This approach is a subset of AI as-a-Service (AIaaS), broadly defined as “cloud-based systems providing on-demand services to organizations and individuals for deploying, developing, training, and managing AI models” (Lins et al., 2021, p. 424). These systems enable organizations to adopt and afford AI technologies without requiring extensive technological expertise (Sundberg and Holmström, 2023; Syed et al., 2025). Moreover, LLMaaS solutions emphasize enhanced data protection and security, offering a viable option for institutions like universities to leverage AI’s advantages while adhering to stringent compliance and security standards (Paavola, 2024).
One proposed advantage of LLMaaS is that organizations can customize various aspects of LLMs to better align with their unique requirements and objectives. In this context, customizations are defined as configurations of parameters at various architectural layers to improve the fit between the system and organization-specific needs (Diaferia et al., 2022). In the broader context of the AI supply chain (Cobbe et al., 2023) such adaptations can occur in a transparent or obscure way for end users (Neumann and Singh, 2025). Table 1 outlines customization options, which encompass customizations of data, models, algorithms, and infrastructure (Diaferia et al., 2022), security and governance adaptations (Chen et al., 2024b), customizations that relate to the design of the user interface and experience, and choices regarding the organizational integration.
Table 1. Examples of customization aspects of AIaaS.
Category Examples References
Model fine-tuning and Configuration Adjust/add system prompts, fine-tuning on proprietary data, hyperparameter configuration, domain-specific knowledge integration, customizable output. (Lins et al., 2021; Diaferia et al., 2022; La Malfa et al., 2023; Hajipour et al., 2023; Parthasarathy et al., 2024; Neumann et al., 2025)
Integration and Interoperability API integration, integration with other tools and systems. (Gan et al., 2023; Lins et al., 2021)
Security and Governance Access control and security, compliance and regulatory requirements, monitoring and evaluation. (Gan et al., 2023; Doosthosseini et al., 2024; Parthasarathy et al., 2024; Iqbal et al., 2024)
User Interface and Experience Incorporation of corporate design, sustainability measures (e.g., token visualization), user profiling and personalization, customizable error handling, multi-language support. (Casati et al., 2019)
Organizational integration Providing context-specific training or workshops, providing context-specific information and use recommendations regarding the model. (Nagarajan et al., 2024)
To date, existing literature has focused on methods and effects of model fine-tuning and configuration, integration and interoperability, as well as security and governance-related consequences of LLMaaS. These aspects are critical because adapting AI systems, for example to ensure data protection or to fine-tune them for internal company data, plays a key role in their secure and human-centered integration into organizations. Accordingly, research and reports on LLMaaS deployment in organizations emphasize primarily technical customizations and their impact on system performance or functionality. For instance, a study on a customized anatomy chatbot showed that tailoring the knowledge base with Retrieval Augmented Generation (RAG) increased factual accuracy (Arun et al., 2024), while reports on the organizational system “FhGenie” highlight aspects such as system architecture and load balancing (Weber et al., 2024). Meanwhile, other customizations pertaining to the user interface of LLMaaS chatbots or their organizational integration – and especially their impact on the end user – receive little attention.
$2.$ Calibration of System Characteristics and User Perceptions: The Role of Cues
Ideally, trust in automated systems should be calibrated, which means that users’ perceived trustworthiness should accurately match the actual capabilities and reliability of the system (Lee and See, 2004). Miscalibration can manifest as either undertrust, leading to system disuse, or overtrust, resulting in system misuse (e. g., failures of monitoring or decision biases) (Parasuraman and Riley, 1997; Wischnewski et al., 2023).
@@
Whether calibrated trust can be achieved depends both on the availability of information about the system and how users perceive and interpret that information (Schlicker et al., 2025). Users’ trust and corresponding use of AI depend on, i. e., “any information element that can be used to make a trust assessment about an agent” (de Visser et al., 2014, p. 253). The TrAM specifies this process, positing that users assess the actual trustworthiness of a system based on available cues, leading to the perceived trustworthiness (Schlicker et al., 2025). The accuracy of this assessment relies on the relevance and availability of cues on the system’s side, as well as the human’s ability to detect and utilize them. On the one hand, actual system characteristics themselves can function as cues, for example information about the training data or the system’s outputs. On the other hand, users may also rely on cues that are less directly tied to a system’s actual qualities, such as interface design or anthropomorphism, which can still strongly influence their perceived trustworthiness (Gong et al., 2024). Such cues may therefore induce unwarranted trust (or distrust) and lead users to perceive systems as more (or less) trustworthy than their actual reliability would justify.
We argue that the concept and relevance of calibration applies not only to trust, but to user perceptions more broadly. For instance, privacy perception calibration can describe the alignment between users’ assumptions about how their data is handled and the system’s actual privacy safeguards. Miscalibrations in this domain have likewise been documented: Windl et al. (2025) showed that misleading cues about browser tracking protection can give users a false sense of privacy. Similarly, hallucination alertness calibration refers to the alignment between users’ cautiousness toward potential incorrect outputs and the system’s actual hallucination tendency. For example, difficulties in detecting biases in LLM outputs have been observed among students in academic use contexts, pointing to a gap between perceived and actual model behavior, a pattern that could similarly occur with hallucinations (Prabhudesai et al., 2025). Accordingly, we propose to conceptualize calibration of user perceptions in the context of LLM use more broadly and understand this as the degree of fit between actual system features (e. g., hallucination tendency, privacy safeguards) and end users’ perceptions of those corresponding system features (e. g., experienced hallucinations or perceived privacy).
@@
Building on the TrAM framework (Schlicker et al., 2025), we focus on the role that customizations of LLMaaS chatbots play as cues that moderate this calibration process between actual and perceived system characteristics. We argue that it is essential to recognize that end users could perceive customizations of LLM chatbots as reflections of the system’s capabilities, goals, and inner workings, even if customizations do not actually affect the functioning of the system. Figure 1 illustrates our proposed interplay between customizations (partially functioning as cues) of the LLMaaS chatbot in relation to actual system characteristics and user perceptions or self-reported behaviors.
@@
Figure 1. System characteristics, customization cues, and user perceptions for the examined university’s customized LLMaaS chatbot. Solid lines represent a positive effect on the associated construct, dotted lines represent a negative effect. Customizations colored in red influence actual system characteristics, while those in blue act as cues expected to affect user perceptions or behavior. Red-and-blue customizations are expected to impact both. Footnotes: ¹Lower temperature may reduce the hallucination tendency by favoring higher-probability tokens and making the output more deterministic (Huang et al., 2025). :
@@ 2.3.1. Background
Trust, as a key determinant of technology adoption (Van Pinxteren et al., 2019; Choudhury and Shamszare, 2023), has been recognized as an important factor for LLM chatbot adoption among academic staff and students (Bhaskar et al., 2024; Tiwari et al., 2024). Research has shown that both insufficient and excessive trust in AI systems can be critical, as a lack of trust may hinder adoption, while overtrust can lead to risks such as reliance on incorrect or potentially harmful outputs (Buçinca et al., 2021; Wagner et al., 2018). This underscores the importance of ensuring that end users place an appropriate – i. e., calibrated – level of trust in LLM chatbots.
Trust is commonly defined as the trustor’s positive expectations towards the trustee, combined with the trustor’s vulnerability and uncertainty within the specific context (Balayn et al., 2025; Lee and See, 2004). Research has shown that trust levels in human-human interaction depend on three perceived qualities of the trustee: ability, integrity, and benevolence (Mayer et al., 1995). In the context of human-automation interaction, these three factors have been reframed as performance, process (e. g., availability, confidentiality, understandability of a system) and purpose (e. g., the developer’s or deployer’s intentions) (Lee and See, 2004). Additional factors have been identified as particularly important for trust in LLMs. These include transparency (Huang et al., 2023; Schwartz et al., 2023), access to a human operator (Nordheim et al., 2019), perceived privacy and robust data security (Huang et al., 2023; Pawłowska-Nowak, 2024).

### v3

@@ Abstract.
As the use of Large Language Model (LLM) chatbots by students and researchers becomes more prevalent, universities are pressed to develop AI strategies. One strategy that many universities pursue is to customize pre-trained LLM-as-a-service (LLMaaS) chatbots. While most studies on LLMaaS chatbots prioritize technical adaptations, these systems are often mainly characterized by user-salient front-end customizations, e. g., interface changes. Yet, no existing studies have examined how users perceive such systems compared to commercial LLM chatbots. In a field study, we investigate how students and employees ($N=526$) at a German university perceive and use their institution’s customized LLMaaS chatbot compared to ChatGPT. Participants using both systems ($n=116$) reported greater trust, higher perceived privacy, and less perceived hallucinations with their university’s customized LLMaaS chatbot compared to ChatGPT. We discuss implications for research on users’ trustworthiness assessment process, and offer guidance for the design and deployment of LLMaaS chatbots.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790622isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Natural language interfacesccs: Security and privacy Privacy protections
@@ 1. Introduction
Large Language Model (LLM) chatbots are becoming increasingly prevalent in educational settings. Current research shows high adoption rates of commercial off-the-shelf LLM chatbots (e. g., ChatGPT) among students and university staff (Zhang et al., 2024; Zhou et al., 2023; von Garrel and Mayer, 2023). Consequently, universities face mounting pressure to develop strategies for AI integration (Kapania et al., 2025). This involves multiple decisions, such as whether and how to provide employees with access to LLMs. To guide the informed and secure use of such systems, many organizations, including universities, corporations and government agencies, are exploring the deployment of LLM-as-a-Service (LLMaaS) as a feasible and cost-effective way to adopt AI in-house (La Malfa et al., 2023; Chkirbene et al., 2024; Odede and Frommholz, 2024). The adoption of such services is driven by several practical considerations, including the opportunity to offer access to all university members while maintaining institutional oversight and alignment with organizational policies (Paavola, 2024). One of the main benefits of LLMaaS is the flexibility they offer in customizing various aspects (Diaferia et al., 2022), such as fine-tuning models or displaying user interfaces in corporate design.
@@
Customizations of LLMaaS solutions play a critical role, as they can enhance model performance within specific organizational contexts and strengthen security and privacy measures (Sundberg and Holmström, 2023; Lewicki et al., 2023). However, even customization choices that have no impact on aspects such as the model or training data, but that are salient to users (e. g., in the form of cues perceivable in the user interface), could lead to significant changes in how the system is perceived and used. The focus of our work is on such front-end customizations, e. g., interface changes, and the role they may play in how users perceive and engage with the system. Psychological research suggests that people draw inferences about a system’s trustworthiness based on indicators available to them, also referred to as cues (de Visser et al., 2014), a mechanism formalized in the Trustworthiness Assessment Model (TrAM) (Schlicker et al., 2025). Real-world incidents show the importance of understanding user perception of LLM chatbots: cases like the data leak by Samsung employees via ChatGPT illustrate how false privacy expectations can lead to the unintended disclosure of sensitive information (Mauran, 2023). Such incidents underscore the importance of understanding user perceptions of differently embedded LLMs, e. g., LLMs offered as a Service vs. commercially available LLMs.
@@
In this work, we examine if customized LLMaaS are perceived and used differently compared to commercially available alternatives. We draw on the TrAM (Schlicker et al., 2025) to theorize on the alignment (or misalignment) between users’ perception of a customized LLMaaS and the system’s actual characteristics, thereby answering the following research question:
  • RQ: How do users’ trust, perception, and usage differ comparing a university’s customized LLMaaS chatbot and a commercially available LLM chatbot?
@@
To examine this research question, we conducted a survey-based field study at a major German university. A diverse sample of students and university employees (N = 526) answered questions about their experiences with the university’s customized LLMaaS chatbot and ChatGPT (the most widely used LLM chatbot among German students (von Garrel and Mayer, 2025)). We measured trust, privacy concerns, cautious behavior toward LLM hallucinations, perceived LLM hallucinations, and sustainability-aware AI usage. Our main analyses focus on a within-subject subsample of participants (n = 116) who regularly used both systems, allowing direct comparison of user perceptions and behaviors across the two platforms.
@@
    We show that participants indicate greater trust, higher perceived privacy and less perceived LLM hallucinations with their university’s customized LLMaaS chatbot compared to ChatGPT.
@@
    Drawing on the TrAM framework (Schlicker et al., 2025), we theorize that customizations of LLMaaS serve as cues that could be the reason for differences in users’ perceptions and behavior toward LLMaaS chatbots compared to commercially available LLM chatbots.
@@ 2.1. LLM (as-a-Service) Chatbots at Universities
LLM chatbots, like ChatGPT, are widely used in academic settings, especially among students (von Garrel and Mayer, 2023; Amani et al., 2023) and researchers (Kapania et al., 2025; Syed et al., 2024; Abdelhafiz et al., 2024). Thus, universities are caught between enabling students and staff to access and use AI in an informed way to leverage the benefits while also protecting data, privacy, rights, and ethical principles. Introducing “their own” LLMs poses a considerable challenge for universities, due to a lack of internal technical abilities and knowledge about the deployment of AI and the high costs of developing and maintaining the required IT infrastructure (Lins et al., 2021).
@@
LLMs-as-a-Service (LLMaaS) provide a practical solution by offering easy access to pre-trained models via cloud providers (Wang et al., 2024; Syed et al., 2025). This approach is a subset of AI-as-a-Service (AIaaS), broadly defined as “cloud-based systems providing on-demand services to organizations and individuals for deploying, developing, training, and managing AI models” (Lins et al., 2021, p. 424). These systems enable organizations to adopt and afford AI technologies without requiring extensive technological expertise (Sundberg and Holmström, 2023; Syed et al., 2025). Moreover, LLMaaS solutions emphasize enhanced data protection and security, offering a viable option for institutions like universities to leverage AI’s advantages while adhering to stringent compliance and security standards (Paavola, 2024).
2.2. Customization of LLMaaS
One proposed advantage of LLMaaS is that organizations can customize various aspects of LLMs to better align with their specific requirements and objectives. Such customizations span different architectural layers – (i) the front-end and (ii) the back-end, which comprise most chatbot applications, (iii) the LLMaaS API, and (iv) the underlying models – to improve the fit between a system and organizational needs (Diaferia et al., 2022). Within the broader AI supply chain (Cobbe et al., 2023), these modifications may be implemented transparently or remain obscure to end users (Neumann and Singh, 2025). Although LLMaaS architectures permit customization across all layers, their implications differ markedly. Back-end, API, or model-level customizations primarily affect technical performance, interoperability, or governance, whereas front-end customizations – the layer directly encountered by users – constitute the most salient and psychologically consequential touchpoints.
Existing research has predominantly examined technical customizations, e. g., model fine-tuning (Arun et al., 2024; Park et al., 2025), integration and interoperability (Weber et al., 2024; Gonzalez Torres and Santa Quintero, 2025), or security and governance considerations (Cai et al., 2025; Weber et al., 2024). In contrast, user-facing customizations remain comparatively underexplored, despite their potential impact on user perceptions, trust, and behavior. These front-end modifications may include transparency features (e. g., explainability panels (Gong et al., 2024)) or visual branding (e. g., logos, color schemes) to align with institutional identity (Casati et al., 2019).
Because branding is frequently intertwined with customization in both practice and scholarly discourse, some conceptual disentangling is necessary. In marketing research, customization is often positioned within broader branding strategies (Yan et al., 2020), yet organizational AI deployment demands a clearer distinction. Branding establishes a recognizable identity through visual or symbolic elements (Bertram, 2016), while customization denotes the wider set of adaptations tailored to institutional needs. Branding-related modifications thus represent only one subset of user-facing customizations. We therefore use the broader term “customization” to capture the full range of user-facing adaptations in LLMaaS deployments, while recognizing branding as one particularly visible – but not exclusive – customization strategy.
In this work, we compare a customized LLMaaS chatbot – primarily modified at the front-end, with only minimal back-end adjustments – to a commercially available alternative (ChatGPT).
2.3. System Characteristics and User Perceptions: The Role of Cues
@@
The aim of this work is to investigate whether users’ perceptions and use of customized LLMaaS chatbots differs from commercially available LLM chatbots. The TrAM model by Schlicker et al. (2025) allows us to theorize on why there might be differences in perceptions and therefore serves as a basis to derive our hypotheses. The TrAM posits that users assess the actual trustworthiness of a system based on available cues, leading to the perceived trustworthiness (Schlicker et al., 2025). Thus, users’ trust and corresponding use of AI depend on, i. e., “any information element that can be used to make a trust assessment about an agent” (de Visser et al., 2014, p. 253). The accuracy of this assessment relies on the relevance and availability of cues on the system’s side, as well as the human’s ability to detect and utilize them. On the one hand, actual system characteristics themselves can function as cues, for example information about the training data or the system’s outputs. On the other hand, users may also rely on cues that are less directly tied to a system’s actual qualities, such as interface design, which can still strongly influence their perceived trustworthiness (Gong et al., 2024). Such cues may therefore induce unwarranted trust (or distrust) and lead users to perceive systems as more (or less) trustworthy than their actual reliability would justify.
The balance or imbalance between perceived vs. warranted trust is also discussed under the umbrella term ’trust calibration’. This concept describes that users’ perceived trustworthiness should accurately match the actual capabilities and reliability of the system (Lee and See, 2004). Miscalibration can manifest as either undertrust, leading to system disuse, or overtrust, resulting in system misuse (e. g., failures of monitoring or decision biases) (Parasuraman and Riley, 1997; Wischnewski et al., 2023). The fit between the actual trustworthiness of a system and its perceived trustworthiness therefore represents one facet of trust calibration (referred to as ‘high accuracy of the trustworthiness assessment’ in the TrAM model; (Schlicker et al., 2025)).
We argue that the concept and relevance of the trustworthiness assessment and the resulting calibration applies not only to trust, but to user perceptions in general. We propose to conceptualize the trustworthiness assessment process with regards to user perceptions in the context of LLM use more broadly and to consider not only users’ trust as an important outcome, but also the perceived privacy, hallucinations as well as sustainable use of the system. Building on the TrAM logic of linking actual system characteristics to perceived system characteristics via cues that are salient to the user, allows us to conceptualize whether there is a fit between actual system features (e. g., hallucination tendency, privacy safeguards) and end users’ perceptions of those corresponding system features (e. g., perceived hallucinations or perceived privacy). We interpret this fit as the degree of calibration of user perceptions with regards to actual system characteristics.
@@
It is essential to recognize that end users could perceive customizations of LLM chatbots as reflections of the system’s capabilities, goals, and inner workings, even if customizations do not actually affect the functioning of the system. Figure 1 illustrates our theoretical reasoning with regards to the expected interplay between customizations (partially functioning as cues) of the LLMaaS chatbot in relation to actual system characteristics and user perceptions or self-reported behaviors. Importantly, this illustrates the basis for the derivation of theoretically-grounded differential hypotheses (see following chapters). These are not mechanisms of action that we explicitly test in an experimental setup. In the following, we elaborate on the theoretical foundations of these expected differences in perception in relation to specific constructs.
@@
Figure 1. Theoretical foundation: System characteristics, customization cues, and user perceptions for the examined university’s customized LLMaaS chatbot. Solid lines represent an expected positive association between the construct, dotted lines represent an expected negative association. Customizations colored in red relate to actual system characteristics, while we expect those in blue to solely relate to perceived characteristics. Red-and-blue customizations are expected to have associations with both. Footnotes: ¹Lower temperature may reduce the hallucination tendency by favoring higher-probability tokens and making the output more deterministic (Huang et al., 2025).
[…]
@@ 2.4.1. Background
Trust, as a key determinant of technology adoption (Van Pinxteren et al., 2019; Choudhury and Shamszare, 2023), has been recognized as an important factor for LLM chatbot adoption (Bhaskar et al., 2024; Tiwari et al., 2024). Trust is commonly defined as the trustor’s positive expectations towards the trustee, combined with the trustor’s vulnerability and uncertainty within the specific context (Balayn et al., 2025; Lee and See, 2004). Research has shown that trust levels in human-human interaction depend on three perceived qualities of the trustee: ability, integrity, and benevolence (Mayer et al., 1995). In the context of human-automation interaction, these three factors have been reframed as performance, process (e. g., availability, confidentiality, understandability of a system) and purpose (e. g., the developer’s or deployer’s intentions) (Lee and See, 2004). Additional factors have been identified as particularly important for trust in LLMs. These include transparency (Huang et al., 2023; Schwartz et al., 2023), access to a human operator (Nordheim et al., 2019), perceived privacy and robust data security (Huang et al., 2023; Pawłowska-Nowak, 2024).